sir i didnt got this question which u taught in video i got {solve for x:sin2x-4cosx=1}

sir  i didnt got this question which u taught in video i got {solve for x:sin2x-4cosx=1}

Grade:11

2 Answers

Akshay
6 years ago
sin2x-4cosx=12sinxcosx - 4cosx = 12cosx(sinx - 2) = 12cosx = 1Cosx =1/2Cosx = cos60°x= π/3 or 60°Sinx - 2 = 1sinx = 3sinx can`t be equal to 3 as sinx lies between [-1,1].
Akshay
6 years ago
sin2x-4cosx=1 ; 2sinxcosx - 4cosx = 1 ; 2cosx(sinx - 2) = 1 ; 2cosx = 1 ; Cosx =1/2 ; Cosx = cos60° ; x= π/3 or 60° ; sinx - 2 = 1 ; sinx = 3 ; sinx can`t be equal to 3 as sinx lies between [-1,1].